OpenStudy (anonymous):

I have a question, in my text it says that "the sum of all the oxidation numbers of all the elements in a neutral molecule is 0. the sum of all the oxidation numbers in a polyatomic ion is equal to the charge on the ion" Does that mean in a polyatomic ion the charge must equal 0? Or is there a different meaning, its really confusing, is it repeating itself? Thanks in advance!

OpenStudy (abb0t):

well, a polyatomic "ion" means there's a charge. But yes, the sum of the oxidation numbers must be neutral. For instance, you have \(H^+\)+ \(OH^-\) you have a +1 and a -1. If you add up the total, your total is zero. By now, you should know from elementary mathematics that -1+1 = 0 Similarly, if you have something like \(Na_2CO_3\), if you separate them Notice you have 2 sodiums per 1 carbonate, which carbonate you know to carry \(2^-\) charge. Now, what if you have a more complex molecule, such as \([MoCl_5]_2\) You have 5 chlorine \(Cl^-\) so 5 \(\times\) -1 = -5, thus Molybdenum must carry an oxidation state of +5. It's a bit different with transition metals since they can carry various charges.

OpenStudy (anonymous):

I get it now. So the the charge of the polyatomic ion and oxidation number are two different numbers. The Polyatomic number is the individual charge of each element. But the oxidation number is the sum of the equation-which must be 0. Am I right?

OpenStudy (abb0t):

oxidation number is the electrical charge that the central atom in a coordination compound would have if all the ligands and electron pairs were removed. usually the oxidation number has the same value as the oxidation state.